Circular No. 8781

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 VZ_{13} (LINEAR) An apparently asteroidal object discovered by the LINEAR survey

(discovery observation tabulated below, taken from MPEC 2006-W03 and MPS 187232) has been found to show a very condensed coma of mag 18.6– 18.7 and diameter 8′′ (with no tail) on R-band CCD images taken by C. W. Hergenrother on Dec. 1.1 and 2.1 UT with the University of Arizona’s 1.54-m Kuiper reflector in bright moonlight.

2006 UT  α_{2000}  δ_{2000}  Mag.
Nov. 13.13467  22^{h}58^{m}19.^{s}72  +42^{o}35′54′′.0  19.9

Parabolic orbital elements from MPEC 2006-X16:

T = 2007 Aug. 11.623 TT  ω = 173.^{o}942
Ω = 66.149 }2000.0
q = 1.02432 AU  i = 134.849

SUPERNOVAE 2006ou, 2006ov, AND 2006ow Two apparent supernovae have been found on unfiltered CCD images:

2006ou by T. Puckett and M. Peoples (cf. IAUC 8779), 2006ov by K. Itagaki (cf. IAUC 8771; via S. Nakano), and 2006ow by T. Boles (cf. IAUC 8716).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ou Nov. 22.41 11^{h}37^{m}13.^{s}10 +15^{o}26′07′′.3 16.0 5′′.9 W, 9′′.6 N
2006ov Nov. 24.86 12 21 55.30 + 4 29 16.7 14.9 5′′.5 E, 51′′ N
2006ow Nov. 26.07 7 35 37.44 +66 24 50.0 17.7 7′′.9 W, 5′′.3 S

Additional magnitudes for 2006ou in UGC 6588: May 22 UT, [19.1; Nov. 25.40, 16.1. SN 2006ou is a type-Ia supernova around a month past maximum (details on CBET 758). Additional unfiltered CCD magnitudes for 2006ov in M61 = NGC 4303 (offset above corrected from CBET 756): May 4, [19.5 (Itagaki); Nov. 25.764, 14.8 (Itagaki); 25.770, 15.0 (K. Kadota, Ageo, Saitama-ken, Japan, 0.25-m reflector; via Nakano). Itagaki reports that nothing is visible at the location of 2006ov on the Digital Sky Survey. SN 2006ov is a type-II supernova, also about a month past maximum (details on CBET 757). Additional magnitudes for 2006ow in UGC 3908: 1989 Dec. 23, [20.5 (Digitized Sky Survey, red); 1997 Mar. 9, [21.0 (DSS, blue); 2005 Nov. 3, [19.5 (Boles); 2006 Jan. 2, [19.5 (Boles); 26.902, 18.0 (Boles). SN 2006ow is a type-Ia supernova past maximum (details on CBET 765).

2006 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green